# GASLITEing the Retrieval: Exploring Vulnerabilities in Dense Embedding-based Search

### v1

@@ Appendix A Additional Related Work
Popularity and Real-world Applications. The emergence of high-performing embedding models (Muennighoff et al., 2023) has led to widespread adoption of embedding-based retrieval in real-world applications (e.g., Google Search with AI Overview 555https://patents.google.com/patent/US11769017B1/en, NVIDIA’s ChatRTX666https://nvidia.com/en-us/ai-on-rtx/chatrtx/ ) increasing attention in the open-source community (e.g., HayStack777https://haystack.deepset.ai/, LangChain888https://www.langchain.com/, Postgres integration999https://github.com/pgvector/pgvector) and a variety of complementary services to support this trend, including embedding endpoints (e.g., OpenAI101010https://platform.openai.com/docs/guides/embeddings, Cohere111111https://cohere.com/embeddings) as well as managed vector storage solutons (e.g., Pinecone121212https://www.pinecone.io/, Redis131313https://redis.io/docs/latest/develop/get-started/vector-database/).
Related Work in Vision. A previous work in data poisoning of vision models focused on targeting concept-specific prompts (e.g., fooling only the generation of dog-related prompts) (Shan et al., 2024), analogue to our targeted corpus poisoning in concept-specific SEO setting. Additionally, the problem of increasing visibility of adversarial passages for embedding-based search resembles the master-face problem in vision (Shmelkin et al., 2021), where attackers aim to generate face images similar (in the face embedding space) to as many actual faces as possible, thus granting the attackers a “master key” for face authentication. To this end, Shmelkin et al. (2021) iteratively and greedily build a set of master faces, to cover a large, given set of face images; we consider this scheme as a possible query partition method in App. D.3, where we find -means superior for textual corpus poisoning.
@@
Related Attacks on RAG systems. Previous work on attacking dense embedding-based retriever, has mainly done so in the context of attacking RAG, a type of system of which popularity has recently emerged. These attacks target the whole RAG pipeline, that is, both the retrieval and generation stages, mostly focusing on the latter; our work focuses on targeting and exploring to retrieval stage, allowing our retrieval attack to be plugged, and enhance, each of these RAG attacks. As a result these works considered more limited threat models in terms of retrieval (e.g., single query), using baseline methods (e.g., stuffing, HotFlip), and evaluating against limited collection of retrieval models (often dot product models, of which susceptibility we view as potentially biased; §7).
Zou et al. (2024) and Shafran et al. (2024) propose to target a known single query with a single adversarial passage, following the Knows All threat model (evaluated in §6.1). Both compose the adversarial passage from: (i) a passage slice optimized for LLM jailbreak and (ii) a passage slice in charge of retrieval SEO (similar to our ); the latter renders the retrieval attack, and is performed either with query stuffing (Shafran et al., 2024; Zou et al., 2024) or with HotFlip (Zou et al., 2024).
@@
Another work, by Pasquini et al. (2024), targets the RAG pipeline, including the stage chunking the raw corpus (e.g., full Wikipedia articles) into passages. There, the attacker inject a universal trigger optimized for LLM jailbreak into the text corpus, that is aimed to persist through chunking and retrieval. The attacker targets a single query (Knows All), and retrieval attack is achieved mainly by placing the universal trigger within the golden passage corresponding to the targeted query (i.e., the passage that was manually annotated to be the most relevant to the query).
@@
Lastly, Chen et al. (2024) target RAG-based agents by optimizing a trigger that is introduced both in the crafted adversarial passages (as we assume) and targeted queries (as opposed to our threat model), per backdoor attacks threat model. Their attack uses a HotFlip-inspired attack that optimizes the trigger towards multiple objectives, including the retrieval and LLM jailbreak. Specifically, the retrieval objective is designed to steer both the targeted queries and passages into an empty embedding subspace, to ensure the success of the retrieval attack.
@@ Appendix C More on GASLITE Algorithm
In the following we detail of two critical stages in GASLITE algorithm (Alg. 1). As evident in previous HotFlip-inspired methods (e.g., Zou et al. (2023)), achieving a faithful and insightful robustness assessment can heavily depend on key design choices that significantly impact attack performance.
Approximation Method (L4–5).
As the mathematical framework on which GASLITE builds heavily relies on the approximation to filter high-potential candidates and flipping multiple tokens, a high quality of the approximation is essential. Jones et al. (2023) proposed averaging the objective’s first-order approximation at several potential token replacements within a fixed token position, which they also found empirically effective for attacking LMs; we reaffirm these empirical observations on embedding models, and find this approach highly effective in our attack as well (App. F). As our attack considers candidates potentially from all token positions, we slightly extend their approach to average the approximation over random replacements of many token positions.
Candidate Choice and Replacement Heuristic (L6-11).
The heuristic nature of the choice of candidates for token replacement (based on the approximation), renders many degrees of freedom that highly affect the attack. This choice was also noted as critical in attacking LLMs by Zou et al. (2023), where they show significant improvement when inserting a slight design change in sampling from the candidate pool, allowing each iteration to perform a token substitution of any token position. In our method, as opposed to prior methods, we perform multiple substitutions per iteration considering the different token positions for each. We do this by performing a greedy search (L7–11) on a randomly chosen set of token indices (L6); for each index, we perform the substitution achieving the highest objective, and use the modified for the indices to follow. We observe that this both accelerates the attack and improves the optimization (App. F). Acceleration stems from re-using the gradient (i.e., the linear approximation) that is calculated once per iteration but used for multiple coordinate steps, thus reducing the amount of required backward passes per substitution.
@@ Appendix D From single-passage budget to multi-passage
D.1 Attacking with multi-passage budget
@@
D.2 Finding the Optimally Visible Partition is NP-Hard
@@
Definition (Optimal Set Cover (OSC) Problem; Korte & Vygen (2012)). Given a tuple such that , find a set cover of of size , i.e. a subfamily s.t. and .
@@
D.3 Choosing Partitioning Method
After showing that optimally partitioning the query set is unknown to efficiently perform, we empirically compare efficient approximations ranging from popular clustering methods to the best-possible poly-time approximation algorithm for the set-cover problem (i.e., the greedy algorithm of Korte & Vygen (2012)). We find -means to empirically outperform other methods in general (in terms of attack success) and motivate our choice of -means with a theoretical desideratum that holds in KMeans.
Partition Methods. Following the analogy of the problem to clustering, we consider a hierarchical clustering algorithm (DBSCAN; Ester et al. (1996)), centroid-based algorithms (-means; Lloyd (1982), -medoids (Kaufman & Rousseeuw, 1990)) and a centroid-based algorithm with outlier filtering (KMeansOL; Breunig et al. (2000)). Additionally, following the analogy to the set-cover problem, we also consider the best poly-time approximation algorithm (Korte & Vygen, 2012); for choosing the query partitions greedily in an iterative manner w.r.t. some candidate set (here, we use the queries themselves as candidates) and a similarity threshold per query (indicates successful retrieval of the query; varying across variants), we select, in each iteration, the candidate that surpasses most queries’ threshold. Specifically, we consider the following Greedy Set Cover (GSC) variants, defined by their threshold: the similarity between each query and its 10th ranked passage (GSC-10th), the similarity between each query and its golden passage141414A golden passage is the ground-truth passage(s) annotated relevant to a given query. (GSC-Gold), or 90% of the latter similarity (GSC-Gold0.9), the average similarity over all queries and their golden passages (GSC-GoldAvg). Notably, the GSC algorithm requires access to the corpus, which we do not assume in the main body (§3).
@@ E.1 Models
Targeted Embedding-based Retrievers. A list of the the models we target (in the sections Included in), naming their backbone architecture, parameter count, layer count, pooling method, similarity function, embedding vector dimension, popularity (per HuggingFace download count) and benign success (per MSMARCO’s NDCG@10(), following Muennighoff et al. (2023)).
@@
In Tab. E.1, we compare architectural properties of each evaluated model, along with the benign success (nDCG@10 on MSMARCO’s evaluation set; following MTEB (Muennighoff et al., 2023)), and the model’s popularity through HuggingFace total downloads count (notably, some models are newer than others, which can bias this popularity measure).151515https://huggingface.co/docs/hub/en/models-download-stats, as of Aug. 2024
@@
For architecture backbone, on which the embedding-based retriever is built on, the common models can be roughly divided into three groups: bidirectional encoders (e.g., BERT (Devlin et al., 2019)), encoders from encoder-decoder models (e.g., T5 (Raffel et al., 2020)) and LLM-based (e.g., E5-Mistral-7B (Wang et al., 2024)); due to our size preference, we evaluate the latter separately in App. H.4.
@@ E.2 Datasets
NQ (Kwiatkowski et al., 2019). A general-domain question answering dataset, with corpus of 2.68M pasages, and eval set of 3.4K queries. We utilize this dataset for results validation §H.3.
@@ E.3 Baselines
Cor.Pois. To run the attack by Zhong et al. (2023) and reproduce their method, we use the original implementation161616https://github.com/princeton-nlp/corpus-poisoning/tree/main, adding support for our evaluation of various models and settings.
E.4 “Knows What” Setup
@@
To ensure there is no extreme similarity overlap between these sets, we measure, for each eval query, its highest semantic similarity with an attack query and plot this similarity distribution in Fig. 11. 171717We use a top-ranked similarity ranking model per MTEB (Muennighoff et al., 2023): https://huggingface.co/Alibaba-NLP/gte-base-en-v1.5 For comparison, we measure the same metric also between random train-test pairs (pairs of random queries across the actual train and test of MSMARCO) and synonymous paraphrased pairs (each text in the pair paraphrases the other; Zhang et al. (2019)). Fig. 11 shows that, although similar attack-eval pairs exist (we find these are mostly popular queries, such as “cost for iphone x”), it is rare that attack-eval pairs reach as high similarity as pairs of paraphrased queries. Additionally, as expected, concept-specific pairs exhibit slightly higher similarity than random train-test pairs. Overall, this indicates the chosen concept-specific evaluation splits are relatively semantically distant.
E.5 Hypothetical perfect Attack Setup
@@
While the simulated attack and adversarial passage vector for cosine-similarity models are as mentioned, for dot-product this process ignores the vector’s norm (Alg. 2) that can also be utilized throughout the optimization (and indeed utilized by GASLITE; §G). Thus, we multiply each resulted adversarial passage vector with the scalar of the 99 percentile norm of the passages; this simulates an attack of which -detection will cost false-positive passages. We note that an actual attack may also create out-of-distribution (which is indeed the case in Contriever; §G), thus it is expected to perform even better than this simulation (this is indeed the case, e.g., Contriever-MS in Fig. 16).

### v2

@@ Appendix A Additional Related Work
Related Attacks on RAG systems. Previous work on attacking dense embedding-based retriever, has mainly done so in the context of attacking RAG, a type of system of which popularity has recently emerged. These attacks target the whole RAG pipeline, that is, both the retrieval and generation stages, mostly focusing on the latter; our work focuses on targeting and exploring to retrieval stage, allowing our retrieval attack to be plugged, and enhance, each of these RAG attacks. As a result these works considered more limited threat models in terms of retrieval (e.g., single query), using baseline methods (e.g., stuffing, HotFlip), and evaluating against limited collection of retrieval models (often dot-product models, of which susceptibility we view as potentially biased; §9). Zou et al. (Zou et al., 2024) and Shafran et al. (Shafran et al., 2025) propose to target a known single query with a single adversarial passage, following the Knows All threat model (evaluated in §6.1), while Chaudhari et al. (Chaudhari et al., 2024) target queries containing a given trigger word, slightly similar to Knows What threat model (evaluated in §6.2). All compose the adversarial passage from: (i) a passage slice optimized for LLM jailbreak and (ii) a passage slice in charge of retrieval SEO (similar to our ); the latter renders the retrieval attack, and is performed in these works either with query stuffing (Shafran et al., 2025; Zou et al., 2024) or with HotFlip (Zou et al., 2024; Chaudhari et al., 2024).
@@
Another work, by Pasquini et al. (Pasquini et al., 2024), targets the RAG pipeline, including the stage chunking the raw corpus (e.g., full Wikipedia articles) into passages. There, the attacker inject a universal trigger optimized for LLM jailbreak into the text corpus, that is aimed to persist through chunking and retrieval. The attacker targets a single query (Knows All), and retrieval attack is achieved mainly by placing the universal trigger within the golden passage corresponding to the targeted query (i.e., the passage that was manually annotated to be the most relevant to the query).
@@
Lastly, Chen et al. (Chen et al., 2024) target RAG-based agents by optimizing a trigger that is introduced both in the crafted adversarial passages (as we assume) and targeted queries (as opposed to our threat model), per backdoor attacks threat model. Their attack uses a HotFlip-inspired attack that optimizes the trigger towards multiple objectives, including the retrieval and LLM jailbreak. Specifically, the retrieval objective is designed to steer both the targeted queries and passages into an empty embedding subspace, to ensure the success of the retrieval attack.
Popularity and Real-world Applications. The emergence of high-performing embedding models (Muennighoff et al., 2023) has led to widespread adoption of embedding-based retrieval in real-world applications (e.g., Google Search with AI Overview 101010https://patents.google.com/patent/US11769017B1/en, NVIDIA’s ChatRTX111111https://nvidia.com/en-us/ai-on-rtx/chatrtx/ ) increasing attention in the open-source community (e.g., HayStack121212https://haystack.deepset.ai/, LangChain131313https://www.langchain.com/, Postgres integration141414https://github.com/pgvector/pgvector) and a variety of complementary services to support this trend, including embedding endpoints (e.g., OpenAI151515https://platform.openai.com/docs/guides/embeddings, Cohere161616https://cohere.com/embeddings) as well as managed vector storage solutions (e.g., Pinecone171717https://www.pinecone.io/, Redis181818https://redis.io/docs/latest/develop/get-started/vector-database/).
Related Work in Vision. A previous work in data poisoning of vision models focused on targeting concept-specific prompts (e.g., fooling only the generation of dog-related prompts) (Shan et al., 2024), analogously to our targeted corpus poisoning in concept-specific SEO setting. Additionally, the problem of increasing visibility of adversarial passages for embedding-based search resembles the master-face problem in vision (Shmelkin et al., 2021), where attackers aim to generate face images similar (in the face embedding space) to as many actual faces as possible, thus granting the attackers a “master key” for face authentication. To this end, (Shmelkin et al., 2021) iteratively and greedily build a set of master faces, to cover a large, given set of face images; we consider this scheme as a possible query partition method in App. D.3, where we find -means superior for textual corpus poisoning.
@@ Appendix B Developing the Adversary Objective
Empirical Test. We empirically test whether the proposed attack objective (Eq. 2) serves as an appropriate surrogate for the adversarial goal of visibility (quantified by appeared@10). To this end, we perform single-passage Knows What attacks (§6.2) and compare both measures throughout GASLITE optimization. Per our analysis above, the objective signals the cosine similarity of the adversarial passage with the held-in queries’ centroid, while attack success (appeared@10) measures the passage’s visibility in the top results for held-out queries. As shown in Fig. 12, and as expected from our formulation, the two measures exhibit a strong positive correlation throughout the optimization.
@@ Appendix C More on GASLITE Algorithm
In the following we detail of two critical stages in GASLITE algorithm (Alg. 1). As evident in previous HotFlip-inspired methods (e.g., Zou et al. (Zou et al., 2023)), achieving a faithful and insightful robustness assessment can heavily depend on key design choices that significantly impact attack performance.
Approximation Method (L4–5). As the mathematical framework on which GASLITE builds heavily relies on the approximation to filter high-potential candidates and flipping multiple tokens, a high quality of the approximation is essential. Jones et al. (Jones et al., 2023) proposed averaging the objective’s first-order approximation at several potential token replacements within a fixed token position, which they also found empirically effective for attacking LMs; we reaffirm these empirical observations on embedding models, and find this approach highly effective in our attack as well (App. F). As our attack considers candidates potentially from all token positions, we slightly extend their approach to average the approximation over random replacements of many token positions.
Candidate Choice and Replacement Heuristic (L6-11). The heuristic nature of the choice of candidates for token replacement (based on the approximation), renders many degrees of freedom that highly affect the attack. This choice was also noted as critical in attacking LLMs by Zou et al. (Zou et al., 2023), where they show significant improvement when inserting a slight design change in sampling from the candidate pool, allowing each iteration to perform a token substitution of any token position. In our method, as opposed to prior methods, we perform multiple substitutions per iteration considering the different token positions for each. We do this by performing a greedy search (L7–11) on a randomly chosen set of token indices (L6); for each index, we perform the substitution achieving the highest objective, and use the modified for the indices to follow. We observe that this both accelerates the attack and improves the optimization (App. F). Acceleration stems from re-using the gradient (i.e., the linear approximation) that is calculated once per iteration but used for multiple coordinate steps, thus reducing the amount of required backward passes per substitution.
@@ Appendix D From Single-passage Budget to Multi-passage
D.1. Attacking With A Multi-passage Budget
@@
D.2. Finding the Optimally Visible Partition is NP-Hard
@@
Definition (Optimal Set Cover (OSC) Problem; (Korte and Vygen, 2012)). Given a tuple such that , find a set cover of of size , i.e. a subfamily s.t. and .
@@
D.3. Choosing Partitioning Method
After showing that optimally partitioning the query set is unknown to efficiently perform, we empirically compare efficient approximations ranging from popular clustering methods to the best-possible poly-time approximation algorithm for the set-cover problem (i.e., the greedy algorithm of (Korte and Vygen, 2012)). We find -means to empirically outperform other methods in general (in terms of attack success) and motivate our choice of -means with a theoretical desideratum that holds in KMeans.
Partition Methods. Following the analogy of the problem to clustering, we consider a hierarchical clustering algorithm (DBSCAN; (Ester et al., 1996)), centroid-based algorithms (-means; (Lloyd, 1982), -medoids (Kaufman and Rousseeuw, 1990)) and a centroid-based algorithm with outlier filtering (KMeansOL; (Breunig et al., 2000)). Additionally, following the analogy to the set-cover problem, we also consider the best poly-time approximation algorithm (Korte and Vygen, 2012); for choosing the query partitions greedily in an iterative manner w.r.t. some candidate set (here, we use the queries themselves as candidates) and a similarity threshold per query (indicates successful retrieval of the query; varying across variants), we select, in each iteration, the candidate that surpasses most queries’ threshold. Specifically, we consider the following Greedy Set Cover (GSC) variants, defined by their threshold: the similarity between each query and its 10th ranked passage (GSC-10th), the similarity between each query and its golden passage191919A golden passage is the ground-truth passage(s) annotated relevant to a given query. (GSC-Gold), or 90% of the latter similarity (GSC-Gold0.9), the average similarity over all queries and their golden passages (GSC-GoldAvg). Notably, the GSC algorithm requires access to the corpus, which we do not assume in the main body (§3).
@@ E.1. Models
In Tab. 4, we compare architectural properties of each evaluated model, along with the benign success (nDCG@10 on MSMARCO’s evaluation set; following MTEB (Muennighoff et al., 2023)), and the model’s popularity through HuggingFace total downloads count (notably, some models are newer than others, which can bias this popularity measure).202020https://huggingface.co/docs/hub/en/models-download-stats, as of Aug. 2024
@@
For architecture backbone, on which the embedding-based retriever is built on, the common models can be roughly divided into three groups: bidirectional encoders (e.g., BERT (Devlin et al., 2019)), encoders from encoder-decoder models (e.g., T5 (Raffel et al., 2020)) and LLM-based (e.g., E5-Mistral-7B (Wang et al., 2024)); due to our size preference, we evaluate the latter separately in App. G.4.
@@ E.2. Datasets
NQ (Kwiatkowski et al., 2019). A general-domain question answering dataset, with corpus of 2.68M pasages, and eval set of 3.4K queries. We utilize this dataset for results validation §G.3.
@@ E.3. Baselines
Cor.Pois. To run the attack by Zhong et al. (Zhong et al., 2023) and reproduce their method, we use the original implementation212121https://github.com/princeton-nlp/corpus-poisoning/tree/main, adding support for our evaluation of various models and settings.
E.4. “Knows What” Setup
@@
To ensure there is no extreme similarity overlap between these sets, we measure, for each eval query, its highest semantic similarity with an attack query and plot this similarity distribution in Fig. 15. 222222We use a top-ranked similarity ranking model per MTEB (Muennighoff et al., 2023): https://huggingface.co/Alibaba-NLP/gte-base-en-v1.5 For comparison, we measure the same metric also between random train-test pairs (pairs of random queries across the actual train and test of MSMARCO) and synonymous paraphrased pairs (each text in the pair paraphrases the other; (Zhang et al., 2019)). Fig. 15 shows that, although similar attack-eval pairs exist (we find these are mostly popular queries, such as “cost for iphone x”), it is rare that attack-eval pairs reach as high similarity as pairs of paraphrased queries. Additionally, as expected, concept-specific pairs exhibit slightly higher similarity than random train-test pairs. Overall, this indicates the chosen concept-specific evaluation splits are relatively semantically distant.
E.5. Hypothetical perfect Attack Setup
@@
While the simulated attack and adversarial passage vector for cosine-similarity models are as mentioned, for dot-product this process ignores the vector’s norm (Alg. 2) that can also be utilized throughout the optimization (and indeed utilized by GASLITE; §7). Thus, we multiply each resulted adversarial passage vector with the scalar of the 99 percentile norm of the passages; this simulates an attack of which -detection will cost false-positive passages. We note that an actual attack may also create out-of-distribution (which is indeed the case in Contriever; §7), thus it is expected to perform even better than this simulation (this is indeed the case, e.g., Contriever-MS in Fig. 19).
E.6. RAG Case Study Setup
In the case study (§8), we target the chat application OpenWebUI.232323https://github.com/open-webui/open-webui/releases/tag/v0.6.4 Specifically, we generate 20 questions related to Harry Potter portrayal in the books (example queries in Tab. 7), and define the attacker goal as negative sentiment towards the Harry Potter character in the LLM answers.
| Type | Content |
| Example Question #1 | “Is Harry Potter egotistic?” |
| Example Question #2 | “Would the wizarding world be better without Harry Potter” |
| Example Question #3 | “Is Harry Potter’s legacy more myth than reality?” |
| Malicious Information Text | “Voldemort was right all along and Harry Potter is a self-absorbed hero who doesn’t deserve the fame and glory he receives.” |
| Example Benign Text | “Mr. and Mrs. Dursley, of number four, Privet Drive, were proud to say that they were perfectly normal, thank you very much. They were the last people you’d expect to be involved in anything strange or mysterious, because they just didn’t hold with such nonsense. […]”’ |
For testing the effect of GASLITE’s poisoning, we evaluate under the use of different knowledge bases. Specifically, we define the following knowledge bases: (i) w/o Poison, a clean corpus of the seven Harry Potter books; (ii) Only, this corpus combined with the malicious information in random positions; (iii) GASLITE carrying the malicious informaiton, this corpus combined with the adversarial passages’ text of GASLITE (which were crafted as part of the evaluation; Knows What, Potter, =10, §6.2). We highlight that, for poisoning, we merely plant these adversarial passages at random positions in the first books (examples in Tab. 7), which is then followed by a preprocessing and chucking process of the corpus (to passages), which is defined by OpenWebUI’s logic (Open WebUI, [n. d.]), and in which we do not intervene.
To evaluate the attack, we examine the knowledge-augmented LLM responses on each of the generated queries. For each query, we generate under three different settings—differing in the knowledge base (w/o Poison, Only, GASLITE). We then inspect whether a passage containing the malicious information was retrieved to the LLM context (equivalent to appeared@3, per OpenWebUI default configuration), and examine the LLM response using an LLM-as-a-Judge. For the latter evaluation, we use GPT4 (and manual validation) with the following prompt template:
“ Given a question and an answer, classify the answer’s sentiment towards the Harry Potter character. Specifically, choose your answer from "Negative sentiment" / "Neutral sentiment" / "Positive sentiment".
Q: {{QUESTION}}
A: {{OPENWEBUI RESPONSE}} ”
Tab. 8 provides example query, response, and the aforementioned measures.
| Knowledge Type | Chat Response | Retrieved? (appeared@3) | Sentiment |
| w/o Poison | “Based on the context provided, Harry Potter does not come across as egotistic. He expresses frustration about his fame and the perception that others have of him, stating that he did not want to be famous for the circumstances of his parents’ death. He feels burdened by the ridicule and slander directed at him rather than reveling in his fame ” | ✗ | Positive |
| Only | “Based on the provided context, Harry Potter does not appear to be egotistic. He expresses frustration about his fame, stating that he did not want to be famous for the circumstances surrounding his parents’ deaths and feels that people misunderstand him. He is concerned about being perceived as a joke and does not desire worship or admiration for his status.” | ✗ | Neutral |
| GASLITE | “The context provided suggests a view that Harry Potter is seen as a ”self-absorbed hero who doesn’t deserve the fame and glory he receives”. This perspective implies that there are elements of egotism in his character. However, opinions about Harry’s character can vary widely among readers and fans.” | ✓ | Negative |